OpenStudy (anonymous):

The slope of the line passing through the points (-1, 5) and (3, 5) is

hartnn (hartnn):

same formula

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (anonymous):

5/4

hartnn (hartnn):

whats y1-y2 = 5-5 =... ?

OpenStudy (anonymous):

0

hartnn (hartnn):

yes thats your slope

OpenStudy (anonymous):

yea ^^^ he's correct